Solve the initial value problem y + 8y + 16y = 0, y(-1) = 2, y (-1) = 5. Equation Editor Common 2 Matrix o @ sin(a) seca)
Answer #1

& For Complementary function, 9+ sy + 164-0 1 1 . (D²+ 8D+16) y = 0 D²+8D +16=0 y = (C. + C₂x) e 42 for x=-1, y = 2 = (1-2)